Limit x tends to 0 (e^x-sinx)/x^2 equals to what please explain in detail

Grade:12

2 Answers

Aradhya gupta
6 years ago
Apply L`hospital rule, differentiate numerator and denominator seperately and get the answer as 1/2=(e^x-cosx)/2x=(e^x+sinx)/2Put x=0=1/2
Aradhya gupta
6 years ago
Limit doesn`t exist.. we can only apply L`hospital when after putting value of x in question we get either 0/0 form or infinity/infinity form.. Here after putting x=0 , we get 1/0 , so L`hospital can`t be applied.. And hence limit doesn`t exist..